# First-Principles Framework for the Prediction of Intersystem Crossing Rates in Spin Defects: The Role of Electron Correlation

### v1

@@ Abstract
Introduction—Optically active spin defects in solids offer a promising platform for advancing quantum technologies [1]. Prototypical systems, such as the negatively charged nitrogen-vacancy (NV-) center in diamond [2, 3] and neutral divacancy centers in silicon carbide [4], have shown potential applications in quantum sensing [5, 6], communication [7, 8, 9, 10], and computation [11, 12] by functioning as spin qubits. First-principles calculations have been instrumental in studying these spin defects [13, 14, 15, 16], in particular their optical spin-polarization cycle, which is critical for the initialization and readout of qubit states [17, 18, 19]. While radiative transitions have been extensively studied using first-principles approaches [20, 21, 22], non-radiative intersystem crossing (ISC) transitions remain less explored. A central challenge lies in accurately accounting for electron correlation effects, which are important for the description of many-body electronic states and their associated spin-orbit coupling (SOC) and electron-phonon (e-ph) interactions [23].
@@
The NV- center, comprising of a nitrogen substitution and an adjacent vacancy, exhibits three localized defect orbitals ($a_{1}$, $e_{x}$, $e_{y}$) within the diamond band gap. In the negatively charged state, four electrons occupy these defect orbitals, resulting in a triplet ground state (${}^{3}\!A_{2}$), a triplet excited state (${}^{3}\!E$), and two singlet states (${}^{1}\!E$ and ${}^{1}\!A_{1}$[24, 25, 26]. The optical spin-polarization cycle (Fig. 1) consists of excitation from the triplet ground to the triplet excited state, followed by ISC to the higher singlet state, with relaxation to a lower singlet state, and finally an ISC process back to the triplet ground state. This cycle enables the initialization and readout of the qubit states, which are defined by the spin sublevels of the ${}^{3}\!A_{2}$ ground state [18].
@@
Here, we focus on ISC transitions from the $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ vibronic levels of ${}^{3}\!E$ to ${}^{1}\!A_{1}$ ($\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ in Fig. 1) and demonstrate the capabilities of our general theoretical and computational framework fully based on first-principles calculations. We combine several methods recently developed to investigate electronic structure properties beyond mean-field approaches. Specifically, we compute SOC using many-body wavefunctions from the quantum defect embedding theory (QDET) [31, 32, 33, 34], which accounts for electron correlation within the active space formed by defect states, while incorporating environmental effects through the screened Coulomb interaction. E-ph interactions are evaluated by computing the vibrational overlap function (VOF) between the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states, using atomic geometries and phonon modes obtained from spin-conserving and spin-flip time-dependent DFT (TDDFT) calculations [35]. This method effectively accounts for the multi-configurational nature of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and has been previously shown to accurately predict the absorption spectrum between singlet states of the NV- center [22]. Additionally, the dynamic Jahn-Teller (DJT) [36] and Herzberg-Teller (HT) effects [37, 38] are incorporated into the VOF calculations. Importantly, our framework relies exclusively on solid-state calculations and is scalable to systems containing hundreds to thousands of atoms, thereby enabling a systematic treatment of finite-size effects for all computed quantities. We first describe the various steps of the computational protocol, and then we compare fluorescence lifetimes with measured values.
Figure 1: Many-body states and optical spin polarization cycle of the NV- center in diamond. The spin-orbit splitting $\lambda_{z}$ represents the energy separation between the sublevels of the ${}^{3}\!E$ state. The intersystem crossing (ISC) rates from vibronic levels $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ to ${}^{1}\!A_{1}$ are labeled as $\Gamma_{A_{1}}$ (red dashed line) and $\Gamma_{E_{1,2}}$ (yellow dashed line). Radiative transitions ${}^{3}\!E\to{}^{3}\!A_{2}$ (blue solid line) and ${}^{1}\!A_{1}\to{}^{1}\!E$ (marron solid line), non-radiative transition ${}^{1}\!A_{1}\to{}^{1}\!E$ (marron dashed line), and ISC ${}^{1}\!E\to{}^{3}\!A_{2}$ (purple dashed line) are shown for completeness.
@@
where $\lambda_{\perp}=1/\sqrt{2}\left\langle{}^{3}\!E(A_{1})\right|\hat{H}_{\mathrm{SO}}\left|{}^{1}\!A_{1}\right\rangle$ is the SOC matrix element between the $A_{1}$ electronic sublevel of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ [25, 26]. Another matrix element, $\lambda_{z}=\left\langle{}^{3}\!E(A_{1})\right|\hat{H}_{\mathrm{SO}}\left|{}^{3}\!E(A_{1})\right\rangle$, determines the splitting of the sublevels in ${}^{3}\!E$. The term $F_{A_{1}}(\Delta)$ is the VOF for the transition from the $\widetilde{A}_{1}$ vibronic level of ${}^{3}\!E$ to the various vibronic levels of ${}^{1}\!A_{1}$. The energy $\Delta$ represents the gap between $\widetilde{A}_{1}$ and the lowest vibronic level of ${}^{1}\!A_{1}$. A detailed derivation of Eq. 1 is provided in Sec. S5 [39]. The ISC rate $\Gamma_{E_{1,2}}$ can be similarly obtained by using the VOF for the $\widetilde{E}_{1,2}$ vibronic level [$F_{E_{1,2}}(\Delta)$].
@@
where $N_{e,\mathcal{A}}$ is the number of electrons in the QDET active space, and $\hat{c}^{\dagger}_{p\sigma}$ ($\hat{c}_{q\sigma^{\prime}}$) are the creation (annihilation) operators for orbital $p$ ($q$) in spin channel $\sigma$ ($\sigma^{\prime}$). The one-body SOC operators entering Eq. 2 are defined as
@@
where $\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}$ and $\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}$ are the non-local (NL) components of the fully relativistic (FR) and the scalar relativistic (SR) pseudopotentials, respectively, for atom $I$ in the supercell [42, 43, 44]. Finite-size effects are addressed by converging the values of $\lambda_{z/\perp}$ with supercell size.
@@
The VOF $F_{A_{1}/E_{1,2}}(\Delta)$ is computed in two steps. First, contributions from $a_{1}$-type phonon modes, $F^{a_{1}}(\Delta)$, are calculated using the Huang-Rhys (HR) theory [50, 20, 21]. This calculation uses the atomic geometries of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ and the phonon modes of ${}^{1}\!A_{1}$, derived from spin-conserving and spin-flip TDDFT calculations [22], which effectively include the multi-configurational nature of these states. Next, the DJT effect of ${}^{3}\!E$, involving $e$-type phonon modes, is incorporated to obtain the final $F_{A_{1}/E_{1,2}}(\Delta)$ [17]. The computed VOFs are extrapolated to the dilute limit using the force constant embedding approach [20, 27]. Equation 1 assumes the validity of the Franck-Condon (FC) principle and treats $\lambda_{\perp}$ as independent of atomic vibrations. Here, we remove this assumption by accounting for the HT effect [37, 38] and compute the derivative of $\lambda_{\perp}$ with respect to configuration coordinates in the calculation of $F^{a_{1}}(\Delta)$.
@@
The energy gap $\Delta$ is approximated as the adiabatic energy difference between ${}^{3}\!E$ and ${}^{1}\!A_{1}$ at their respective equilibrium atomic geometries, assuming approximate cancellation of their zero-point energies. The vertical energy gap between ${}^{3}\!E$ and ${}^{1}\!A_{1}$ is computed using QDET, while their FC shifts (reorganization energies) are obtained from TDDFT [22, 35]. Further details on our computational framework are in Secs. S1 and S2 [39].
@@
Results—We first compute the SOC parameter, $\lambda_{z/\perp}$, using many-body wavefunctions derived from group theory, using the minimal basis set of defect orbitals $a_{1}$, $e_{x}$, and $e_{y}$ (see Tables S1 and S2 [39]). The resulting expressions are: $\lambda_{z}=|\langle e_{x}|\hat{\mathbf{z}}|e_{y}\rangle|/2$ and $\lambda_{\perp}=|\langle a_{1}|\hat{\mathbf{z}}|e_{x}\rangle|/2$. Results obtained from these expressions are denoted as $\lambda_{z/\perp}^{\mathrm{DFT}}$. Figure 2(a) shows the strong dependence of $\lambda_{z/\perp}^{\mathrm{DFT}}$ on supercell size, attributed to the overlap of defect orbitals with periodic images in small supercells. Extrapolating to the dilute limit via exponential decay yields $\lambda_{z,0}^{\mathrm{DFT}}=37.5$ GHz and $\lambda_{\perp,0}^{\mathrm{DFT}}=70.6$ GHz.
@@
To properly account for electron correlation effects beyond the group theory analysis, we compute $\lambda_{z/\perp}$ using many-body wavefunctions from QDET calculations, denoted as $\lambda_{z/\perp}^{\mathrm{QDET}}$. Figure 2(b) presents $\lambda_{z/\perp}^{\mathrm{QDET}}$ for a 511-atom supercell, highlighting the significant influence of electron correlation. While QDET wavefunctions resemble those from the group theory analysis, they do contain additional configurations that introduce small yet important corrections, leading to a notable decrease in $\lambda_{z/\perp}^{\mathrm{QDET}}$ compared to $\lambda_{z/\perp}^{\mathrm{DFT}}$ (details in Secs. S3 and S4 [39]). By accounting for the weak dependence of $\lambda_{z/\perp}^{\mathrm{QDET}}$ on the active space choice and applying a finite-size correction (assumed to be the same as for $\lambda_{z/\perp}^{\mathrm{DFT}}$), we determine the final SOC parameters as $\lambda_{z,0}^{\mathrm{QDET}}=23.1\pm 4.3$ GHz and $\lambda_{\perp,0}^{\mathrm{QDET}}=29.4\pm 2.8$ GHz. These values are used in the ISC rate calculations presented below.
@@
In Fig. 2(c), we compare our computed $\lambda_{z/\perp}$ with previous theoretical and experimental results. The quantity $p\lambda_{z}$ represents the measured fine structure splitting of ${}^{3}\!E$, where $p$ accounts for the Ham reduction factor due to the DJT effect. Using $p=0.304$ from Ref. [17], we obtain $p\lambda_{z,0}^{\mathrm{QDET}}=7.0\pm 1.3$ GHz, in good agreement with the measured values: $5.33\pm 0.03$ GHz [46, 47], whereas $p\lambda_{z,0}^{\mathrm{DFT}}$ overestimates experiment by a factor of two. Our computed $\lambda_{\perp,0}^{\mathrm{QDET}}$ also aligns well with the experimental estimate based on $\lambda_{\perp}/\lambda_{z}=1.2\pm 0.2$ relation [48, 49], while $\lambda_{\perp,0}^{\mathrm{DFT}}$ is approximately three times as large. We note that previous quantum chemistry CASSCF calculations [29] significantly underestimated $\lambda_{\perp}$, likely due to an inaccurate description of the NV- center’s many-body wavefunctions with a small C33H36N- cluster.
@@
As mentioned above, in our calculations of VOFs for the $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ ISC transitions, we first considered contributions from $a_{1}$-type phonon modes. The computed $F^{a_{1}}(\Delta)$ is shown in Fig. 3(a). Previous studies often derived $F^{a_{1}}(\Delta)$ from constrained-occupations DFT ($\Delta$SCF), which approximates ${}^{3}\!E$ and ${}^{1}\!A_{1}$ as single Slater-determinant states and inadequately account for electron correlation effects [17]; alternatively, $F^{a_{1}}(\Delta)$ was approximated by the VOF for the ${}^{3}\!E\to{}^{3}\!A_{2}$ photoluminescence (PL) transition [48, 49, 29]. Here, we compute $F^{a_{1}}(\Delta)$ using the HR theory [50], based on equilibrium atomic geometries of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and $a_{1}$-type phonon modes of ${}^{1}\!A_{1}$, obtained from spin-conserving and spin-flip TDDFT calculations. A scaling factor is applied to adjust the HR factors, correcting for underestimated atomic displacements due to the use of the PBE functional [27, 22] (see Sec. S6 [39]). We find notable differences between $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ and ${}^{3}\!E\to{}^{3}\!A_{2}$ transitions [Fig. 3(a)]. Specifically, in the range $\Delta\in[0.3,0.4]$ eV, $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{3}\!A_{2}$ exceeds that of ${}^{3}\!E\to{}^{1}\!A_{1}$ by a factor of 1.5–1.9. Furthermore, the ${}^{3}\!E\to{}^{3}\!A_{2}$ PL lineshape overestimates $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ by a factor of 1.9–3.1 due to contributions from $e$-type phonon modes.
@@
The VOF computed using the HR theory assumes the validity of the FC principle, where $\lambda_{\perp}$ is treated as constant between the atomic geometries of the initial and final states of the ISC transition. However, we find that $\lambda_{\perp}$ varies by 8% between the ${}^{1}\!A_{1}$ and ${}^{3}E$ equilibrium atomic geometries (see Sec. S7 [39]). This variation is accounted for by including the HT effect [37, 38] in our calculations. Using the ${}^{1}\!A_{1}$ equilibrium atomic geometry as the reference, we compute the Franck-Condon Herzberg-Teller (FCHT) and HT VOFs, shown in Fig. 3(b). Our results reveal that while the HT term contributes only 1%, the FCHT term reduces the FC term intensity by 19%. These findings indicate that the HT effect has a minor influence on the VOF in the NV- center. However, our calculations demonstrate the capability of the framework developed here to assess vibronic coupling effects in SOC calculations [51].
@@
Finally, using an effective Hamiltonian describing the interaction between ${}^{3}\!E$ electronic states and $e$-type phonon modes [17], we incorporate the DJT effect into the VOF calculations for $\Gamma_{A_{1}}$ and the $\Gamma_{E_{1,2}}$, yielding $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ [Fig. 3(c)]. The DJT effect mixes the $A_{1}$, $A_{2}$, and $E_{1,2}$ electronic levels of ${}^{3}\!E$, giving rise to the vibronic states $\widetilde{A}_{1}$, $\widetilde{A}_{2}$, and $\widetilde{E}_{1,2}$ through coupling with $e$-type phonons. This coupling reduces the intensity of $F_{A_{1}}(\Delta)$ relative to the bare $F^{a_{1}}(\Delta)$ and generates a non-zero $F_{E_{1,2}}(\Delta)$, thereby enabling the $\Gamma_{E_{1,2}}$ ISC transition. The onset of $F_{E_{1,2}}(\Delta)$ at 78 meV confirm its $e$-type phonon-assisted nature, consistent with the model of Ref. [48]. Further details are provided in Sec. S5 [39].
@@
We now combine the calculations of the SOC parameter $\lambda_{\perp,0}^{\mathrm{QDET}}$ and the VOFs $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ to compute the ISC rates $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ as functions of the energy gap $\Delta$. Results for $\Gamma_{A_{1}}$ are shown in Fig. 4(a), while those for $\Gamma_{E_{1,2}}$ are in Fig. S10 [39]. Comparing the computed $\Gamma_{A_{1}}$ with experimental data [52, 48] yields an estimated $\Delta$ range of $[0.334,0.389]$ eV. Similarly, the comparison of $\Gamma_{E_{1,2}}$ with experimental results [52, 48] gives $[0.314,0.375]$ eV. The overlap of these two ranges provides a refined estimate of $\Delta\in[0.334,0.375]$ eV, offering a reliable prediction for the ${}^{3}\!E-{}^{1}\!A_{1}$ energy gap. The ISC rates computed using $\lambda_{\perp,0}^{\mathrm{DFT}}$, also shown in Fig. 4, overestimate experiments by roughly a factor of 6. The inverse dependence of $\Gamma_{A_{1}/E_{1,2}}$ on $\Delta$ suggests that the NV- center’s ISC rates, and hence the optical spin-readout contrast, could be enhanced by decreasing $\Delta$, potentially through strain engineering [53, 54].

### v2

@@ Abstract
Introduction—Optically active spin defects in solids offer a promising platform for advancing quantum technologies [1]. Prototypical systems, such as the negatively charged nitrogen-vacancy (NV-) center in diamond [2, 3] and neutral divacancy centers in silicon carbide [4], have shown potential applications in quantum sensing [5, 6], communication [7, 8, 9, 10], and computation [11, 12] by functioning as spin qubits. First-principles calculations have been instrumental in studying these spin defects [13, 14, 15, 16], in particular their optical spin-polarization cycle, which is critical for the initialization and readout of qubit states [17, 18, 19]. While radiative transitions have been extensively studied using first-principles approaches [20, 21, 22], nonradiative intersystem crossing (ISC) transitions remain less explored. A central challenge lies in accurately accounting for electron correlation effects, which are important for the description of many-body electronic states and their associated spin-orbit coupling (SOC) and electron-phonon (e-ph) interactions [23].
@@
The NV- center, comprising a nitrogen substitution and an adjacent vacancy, exhibits three localized defect orbitals ($a_{1}$, $e_{x}$, $e_{y}$) within the diamond band gap. In the negatively charged state, four electrons occupy these defect orbitals, resulting in a triplet ground state (${}^{3}\!A_{2}$), a triplet excited state (${}^{3}\!E$), and two singlet states (${}^{1}\!E$ and ${}^{1}\!A_{1}$[24, 25, 26]. The optical spin-polarization cycle (Fig. 1) consists of excitation from the triplet ground to the triplet excited state, followed by ISC to the higher singlet state, with relaxation to a lower singlet state, and finally an ISC process back to the triplet ground state. This cycle enables the initialization and readout of the qubit states, which are defined by the spin sublevels of the ${}^{3}\!A_{2}$ ground state [18].
@@
Here, we focus on ISC transitions from the $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ vibronic levels of ${}^{3}\!E$ to ${}^{1}\!A_{1}$ ($\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ in Fig. 1) and demonstrate the capabilities of our general theoretical and computational framework fully based on first-principles calculations. We combine several methods recently developed to investigate electronic structure properties beyond mean-field approaches. Specifically, we compute SOC using many-body wave functions from the quantum defect embedding theory (QDET) [31, 32, 33, 34], which accounts for electron correlation within the active space formed by defect states, while incorporating environmental effects through the screened Coulomb interaction. E-ph interactions are evaluated by computing the vibrational overlap function (VOF) between the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states, using atomic geometries and phonon modes obtained from spin-conserving and spin-flip time-dependent DFT (TDDFT) calculations [35]. This method effectively accounts for the multiconfigurational nature of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and has been previously shown to accurately predict the absorption spectrum between singlet states of the NV- center [22]. Additionally, the dynamic Jahn-Teller (DJT) [36] and Herzberg-Teller (HT) effects [37, 38] are incorporated into the VOF calculations. Importantly, our framework relies exclusively on solid-state calculations and is scalable to systems containing hundreds to thousands of atoms, thereby enabling a systematic treatment of finite-size effects for all computed quantities. We first describe the various steps of the computational protocol, and then we compare fluorescence lifetimes with measured values.
Figure 1: Many-body states and optical spin-polarization cycle of the NV- center in diamond. The spin-orbit splitting $\lambda_{z}$ represents the energy separation between the sublevels of the ${}^{3}\!E$ state. The intersystem crossing (ISC) rates from vibronic levels $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ to ${}^{1}\!A_{1}$ are labeled as $\Gamma_{A_{1}}$ (red dashed line) and $\Gamma_{E_{1,2}}$ (yellow dashed line). Radiative transitions ${}^{3}\!E\to{}^{3}\!A_{2}$ (blue solid line) and ${}^{1}\!A_{1}\to{}^{1}\!E$ (maroon solid line), nonradiative transition ${}^{1}\!A_{1}\to{}^{1}\!E$ (maroon dashed line), and ISC ${}^{1}\!E\to{}^{3}\!A_{2}$ (purple dashed line) are shown for completeness.
@@
where $\lambda_{\perp}=1/\sqrt{2}\left\langle{}^{3}\!E(A_{1})\right|\hat{H}_{\mathrm{SO}}\left|{}^{1}\!A_{1}\right\rangle$ is the SOC matrix element between the $A_{1}$ electronic sublevel of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ [25, 26]. Another matrix element, $\lambda_{z}=\left\langle{}^{3}\!E(A_{1})\right|\hat{H}_{\mathrm{SO}}\left|{}^{3}\!E(A_{1})\right\rangle$, determines the splitting of the sublevels in ${}^{3}\!E$. The term $F_{A_{1}}(\Delta)$ is the VOF for the transition from the $\widetilde{A}_{1}$ vibronic level of ${}^{3}\!E$ to the various vibronic levels of ${}^{1}\!A_{1}$. The energy $\Delta$ represents the gap between $\widetilde{A}_{1}$ and the lowest vibronic level of ${}^{1}\!A_{1}$. A detailed derivation of Eq. (1) is provided in Sec. S5 of Supplemental Material (SM) [39]. The ISC rate $\Gamma_{E_{1,2}}$ can be similarly obtained by using the VOF for the $\widetilde{E}_{1,2}$ vibronic level [$F_{E_{1,2}}(\Delta)$].
@@
where $N_{e,\mathcal{A}}$ is the number of electrons in the QDET active space, and $\hat{c}^{\dagger}_{p\sigma}$ ($\hat{c}_{q\sigma^{\prime}}$) are the creation (annihilation) operators for orbital $p$ ($q$) in spin channel $\sigma$ ($\sigma^{\prime}$). The one-body SOC operators entering Eq. (2) are defined as
@@
where $\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}$ and $\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}$ are the nonlocal (NL) components of the fully relativistic (FR) and the scalar relativistic (SR) pseudopotentials, respectively, for atom $I$ in the supercell [42, 43, 44]. Finite-size effects are addressed by converging the values of $\lambda_{z/\perp}$ with supercell size.
@@
The VOF $F_{A_{1}/E_{1,2}}(\Delta)$ is computed in two steps. First, contributions from $a_{1}$-type phonon modes, $F^{a_{1}}(\Delta)$, are calculated using the Huang-Rhys (HR) theory [50, 20, 21]. This calculation uses the atomic geometries of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ and the phonon modes of ${}^{1}\!A_{1}$, derived from spin-conserving and spin-flip TDDFT calculations [22], which effectively include the multiconfigurational nature of these states. Next, the DJT effect of ${}^{3}\!E$, involving $e$-type phonon modes, is incorporated to obtain the final $F_{A_{1}/E_{1,2}}(\Delta)$ [17]. The computed VOFs are extrapolated to the dilute limit using the force constant embedding approach [20, 27]. Equation (1) assumes the validity of the Franck-Condon (FC) principle and treats $\lambda_{\perp}$ as independent of atomic vibrations. Here, we remove this assumption by accounting for the HT effect [37, 38] and compute the derivative of $\lambda_{\perp}$ with respect to configuration coordinates in the calculation of $F^{a_{1}}(\Delta)$.
@@
The energy gap $\Delta$ is approximated as the adiabatic energy difference between ${}^{3}\!E$ and ${}^{1}\!A_{1}$ at their respective equilibrium atomic geometries, assuming approximate cancellation of their zero-point energies. The vertical energy gap between ${}^{3}\!E$ and ${}^{1}\!A_{1}$ is computed using QDET, while their FC shifts (reorganization energies) are obtained from TDDFT [22, 35]. Further details on our computational framework are in Secs. S1 and S2 of SM [39]. Although not explored here, this computational framework is readily applicable to ISC transitions from ${}^{1}\!E$ to ${}^{3}\!A_{2}$ ($t_{\pm}$ and $t_{z}$ in Fig. 1) by substituting the SOC matrix element, the VOF, and the energy gap with those between ${}^{1}\!E$ and ${}^{3}\!A_{2}$.
@@
Results—We first compute the SOC parameter, $\lambda_{z/\perp}$, using many-body wave functions derived from group theory, using the minimal basis set of defect orbitals $a_{1}$, $e_{x}$, and $e_{y}$ (see Tables S1 and S2 in SM [39]). The resulting expressions are – $\lambda_{z}=|\langle e_{x}|\hat{\mathbf{z}}|e_{y}\rangle|/2$ and $\lambda_{\perp}=|\langle a_{1}|\hat{\mathbf{z}}|e_{x}\rangle|/2$. Results obtained from these expressions are denoted as $\lambda_{z/\perp}^{\mathrm{DFT}}$. Figure 2(a) shows the strong dependence of $\lambda_{z/\perp}^{\mathrm{DFT}}$ on supercell size, attributed to the overlap of defect orbitals with periodic images in small supercells. Extrapolating to the dilute limit via exponential decay yields $\lambda_{z,0}^{\mathrm{DFT}}=37.5$ GHz and $\lambda_{\perp,0}^{\mathrm{DFT}}=70.6$ GHz.
@@
To properly account for electron correlation effects beyond the group theory analysis, we compute $\lambda_{z/\perp}$ using many-body wave functions from QDET calculations, denoted as $\lambda_{z/\perp}^{\mathrm{QDET}}$. Figure 2(b) presents $\lambda_{z/\perp}^{\mathrm{QDET}}$ for a 511-atom supercell, highlighting the significant influence of electron correlation. While QDET wave functions resemble those from the group theory analysis, they do contain additional configurations that introduce small yet important corrections, leading to a notable decrease in $\lambda_{z/\perp}^{\mathrm{QDET}}$ compared to $\lambda_{z/\perp}^{\mathrm{DFT}}$ (details in Secs. S3 and S4 [39] of SM). By accounting for the weak dependence of $\lambda_{z/\perp}^{\mathrm{QDET}}$ on the active space choice and applying a finite-size correction (assumed to be the same as for $\lambda_{z/\perp}^{\mathrm{DFT}}$), we determine the final SOC parameters as $\lambda_{z,0}^{\mathrm{QDET}}=23.1\pm 4.3$ GHz and $\lambda_{\perp,0}^{\mathrm{QDET}}=29.4\pm 2.8$ GHz. These values are used in the ISC rate calculations presented below.
@@
In Fig. 2(c), we compare our computed $\lambda_{z/\perp}$ with previous theoretical and experimental results. The quantity $p\lambda_{z}$ represents the measured fine structure splitting of ${}^{3}\!E$, where $p$ accounts for the Ham reduction factor due to the DJT effect. Using $p=0.304$ from Ref. [17], we obtain $p\lambda_{z,0}^{\mathrm{QDET}}=7.0\pm 1.3$ GHz, in good agreement with the measured values, $5.33\pm 0.03$ GHz [46, 47], whereas $p\lambda_{z,0}^{\mathrm{DFT}}$ overestimates experiment by a factor of 2. Our computed $\lambda_{\perp,0}^{\mathrm{QDET}}$ also aligns well with the experimental estimate based on $\lambda_{\perp}/\lambda_{z}=1.2\pm 0.2$ relation [48, 49], while $\lambda_{\perp,0}^{\mathrm{DFT}}$ is approximately 3 times as large. We note that previous complete active space self-consistent field (CASSCF) calculations [29] significantly underestimated $\lambda_{\perp}$, likely due to an inaccurate description of the NV- center’s many-body wave functions with a small C33H36N- cluster.
@@
As mentioned above, in our calculations of VOFs for the $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ ISC transitions, we first considered contributions from $a_{1}$-type phonon modes. The computed $F^{a_{1}}(\Delta)$ is shown in Fig. 3(a). Previous studies often derived $F^{a_{1}}(\Delta)$ from constrained-occupations DFT ($\Delta$SCF), which approximates ${}^{3}\!E$ and ${}^{1}\!A_{1}$ as single Slater-determinant states and inadequately account for electron correlation effects [17]; alternatively, $F^{a_{1}}(\Delta)$ was approximated by the VOF for the ${}^{3}\!E\to{}^{3}\!A_{2}$ photoluminescence (PL) transition [48, 49, 29]. Here, we compute $F^{a_{1}}(\Delta)$ using the HR theory [50], based on equilibrium atomic geometries of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and $a_{1}$-type phonon modes of ${}^{1}\!A_{1}$, obtained from spin-conserving and spin-flip TDDFT calculations. A scaling factor is applied to adjust the HR factors, correcting for underestimated atomic displacements due to the use of the Perdew-Burke-Emzerhof (PBE) functional [27, 22] (see Sec. S6 of SM [39]). We find notable differences between $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ and ${}^{3}\!E\to{}^{3}\!A_{2}$ transitions [Fig. 3(a)]. Specifically, in the range $\Delta\in[0.3,0.4]$ eV, $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{3}\!A_{2}$ exceeds that of ${}^{3}\!E\to{}^{1}\!A_{1}$ by a factor of 1.5–1.9. Furthermore, the ${}^{3}\!E\to{}^{3}\!A_{2}$ PL line shape overestimates $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ by a factor of 1.9–3.1 due to contributions from $e$-type phonon modes.
@@
The VOF computed using the HR theory assumes the validity of the FC principle, where $\lambda_{\perp}$ is treated as constant between the atomic geometries of the initial and final states of the ISC transition. However, we find that $\lambda_{\perp}$ varies by 8% between the ${}^{1}\!A_{1}$ and ${}^{3}E$ equilibrium atomic geometries (see Sec. S7 of SM [39]). This variation is accounted for by including the HT effect [37, 38] in our calculations. Using the ${}^{1}\!A_{1}$ equilibrium atomic geometry as the reference, we compute the Franck-Condon Herzberg-Teller (FCHT) and HT VOFs, shown in Fig. 3(b). Our results reveal that while the HT term contributes only 1%, the FCHT term reduces the FC term intensity by 19%. These findings indicate that the HT effect has a minor influence on the VOF in the NV- center. However, our calculations demonstrate the capability of the framework developed here to assess vibronic coupling effects in SOC calculations [51].
@@
Finally, using an effective Hamiltonian describing the interaction between ${}^{3}\!E$ electronic states and $e$-type phonon modes [17], we incorporate the DJT effect into the VOF calculations for $\Gamma_{A_{1}}$ and the $\Gamma_{E_{1,2}}$, yielding $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ [Fig. 3(c)]. The DJT effect mixes the $A_{1}$, $A_{2}$, and $E_{1,2}$ electronic levels of ${}^{3}\!E$, giving rise to the vibronic states $\widetilde{A}_{1}$, $\widetilde{A}_{2}$, and $\widetilde{E}_{1,2}$ through coupling with $e$-type phonons. This coupling reduces the intensity of $F_{A_{1}}(\Delta)$ relative to the bare $F^{a_{1}}(\Delta)$ and generates a nonzero $F_{E_{1,2}}(\Delta)$, thereby enabling the $\Gamma_{E_{1,2}}$ ISC transition. The onset of $F_{E_{1,2}}(\Delta)$ at 78 meV confirm its $e$-type phonon-assisted nature, consistent with the model of Ref. [48]. Further details are provided in Sec. S5 of SM [39].
@@
We now combine the calculations of the SOC parameter $\lambda_{\perp,0}^{\mathrm{QDET}}$ and the VOFs $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ to compute the ISC rates $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ as functions of the energy gap $\Delta$. Results for $\Gamma_{A_{1}}$ are shown in Fig. 4(a), while those for $\Gamma_{E_{1,2}}$ are in Fig. S10 of SM [39]. Comparing the computed $\Gamma_{A_{1}}$ with experimental data [52, 48] yields an estimated $\Delta$ range of $[0.334,0.389]$ eV. Similarly, the comparison of $\Gamma_{E_{1,2}}$ with experimental results [52, 48] gives $[0.314,0.375]$ eV. The overlap of these two ranges provides a refined estimate of $\Delta\in[0.334,0.375]$ eV, offering a reliable prediction for the ${}^{3}\!E-{}^{1}\!A_{1}$ energy gap. The ISC rates computed using $\lambda_{\perp,0}^{\mathrm{DFT}}$, also shown in Fig. 4, overestimate experiments by roughly a factor of 6. The inverse dependence of $\Gamma_{A_{1}/E_{1,2}}$ on $\Delta$ suggests that the NV- center’s ISC rates, and hence the optical spin-readout contrast, could be enhanced by decreasing $\Delta$, potentially through strain engineering [53, 54].